# Security Weaknesses of Copilot Generated Code in GitHub

### v1

@@ 3.1. Research Goal and Questions
This empirical study aims to understand the potential security weaknesses in Copilot generated code. We first collected code snippets generated by Copilot from GitHub projects as the data source for our research. It should be noted that it is not possible to access all the code generated by Copilot in GitHub projects, as there is no direct way to identify if part of a file was generated by Copilot (i.e., source files do not contain any signatures to indicate if Copilot generates the code). However, we can identify many code snippets by searching through the repository description and the comments provided in the code (details provided in Section 3.2.2). We then analyzed the identified snippets for security weaknesses. We aim to help developers and researchers understand common security weaknesses when using Copilot without focusing on whether the code generation aspects are used correctly.
We conducted this empirical study following the guidelines proposed in (Easterbrook et al., 2008). The RQs, their rationale, and the research process of this study (see Fig. 1) are detailed in the subsections below.
@@
Rationale: Copilot may produce code suggestions that developers accept but these suggestions may include security weaknesses that could potentially make the program vulnerable. The answer to RQ1 helps to understand the frequency of security weaknesses developers encounter when using Copilot in production.
@@
Rationale: Copilot generated code may contain security weaknesses (Pearce et al., 2022), and developers should conduct a rigorous security review before accepting the code generated by Copilot. As clarified in the documentation of GitHub Copilot “the users of Copilot are responsible for ensuring the security and quality of their code (GitHub, [n. d.])”. The answer to RQ2 can help developers better understand possible security weaknesses in the code generated by Copilot, thereby enabling them to prevent and fix these weaknesses more effectively.
RQ3. How many security weaknesses belong to the MITRE CWE Top-25?
Rationale: The MITRE list contains the top 25 most dangerous security weaknesses. The answer to RQ3 can help developers understand whether the code generated by Copilot contains widely recognized types of security weaknesses and Copilot’s ability to handle these recent and common weaknesses.
@@ 3.2. Data Collection and Filtering
We chose GitHub as the primary data source for answering our RQs. GitHub is widely used by developers. As the world’s largest code hosting platform, GitHub contains millions of public code repositories and offers access to a large number of code resources, allowing us to cover multiple programming languages and project types in our study. We used code snippets generated by Copilot on GitHub as our research object to analyze the relevant security weaknesses of Copilot. Our scripts and dataset are provided online in our replication package (Fu et al., 2023).
@@ 3.2.1. Code Snippets Collection
Step 1. To collect code snippets generated by Copilot, we first conducted a pilot search to formulate our search keywords. First, we used “GitHub Copilot” and “Copilot” as our search keywords. As expected, we found that the term “Copilot” not only refers to the code generation tool launched by GitHub but also to tools in the aviation or telemetry fields. Therefore, using the keyword “Copilot” solely may return irrelevant content that may not be related to the use of the tool. On the other hand, using “GitHub Copilot” as a search keyword can exclude content unrelated to the code generation tool Copilot and narrow the search scope, which is what we have used to locate the code snippets.
However, even with this basic search keyword, we still need to carefully filter the search results to ensure they are truly related to GitHub Copilot. Although using “GitHub Copilot” as a search keyword increases the relevance of the results to Copilot, these results are not necessarily the code snippets generated by Copilot. It should be noted that many code snippets containing the “GitHub Copilot” keyword in the search results display GitHub Copilot as text. Developers may use them to describe their experience using Copilot to generate code or showcase information related to Copilot. These code snippets are not what we need because they do not directly relate to the code generated by Copilot. Our target is code generated by Copilot, not code snippets containing the keyword “Copilot”.
Our observations from the pilot search showed that using keywords such as “by GitHub Copilot”, “use GitHub Copilot”, and “with GitHub Copilot” can improve the accuracy of the search results. These keywords enable us to focus more on the code generated using Copilot rather than code snippets that contain other content related to Copilot. In addition, since our goal is to use automated analysis tools to perform security scans on the collected code snippets, we further limited the types of code snippets during the search to Python, JavaScript, Java, C++, C#, and Go. These are the mainstream languages supported by Copilot, and also the languages supported by CodeQL. We collected the Code parts from these search results. Considering that some projects declare using GitHub Copilot generated code in their README files or project description provided in GitHub, we decided to retain the results from the Repository label in the search results. Fig. 2 shows an example of our search process.
Table 1 reports the search terms we used and the number of search results obtained from GitHub. In this step, we collected a total of 8,004 results, of which 7,749 were from the Code label, and 255 were from the Repository label. The same search result may contain multiple keywords, meaning there are duplicate projects in the collected data. After removing duplicate projects, we obtained a total of 4249 search results, of which 4081 were from the Code label, and 168 were from the Repository label. Table 2 shows the number of different language types of search results we obtained from the Code label.
@@
1
“By GitHub Copilot”
2549
54
2
“Use GitHub Copilot”
1822
77
3
“With GitHub Copilot”
3378
127
Total 7749 255
Table 2. Search Results from GitHub
# Language # Code results
L1
Python
784
L2
JavaScript
863
L3
Java
641
L4
C++
437
L5
C#
386
L6
Go
970
Total 4081
@@ 3.2.2. Filtering Code Snippets
Step 2. After obtaining the results from the keyword searches, we further filtered them by not only considering the accuracy of the keywords but also investigating the project’s documentation, code comments, and other metadata in the search results to determine whether they were generated by GitHub Copilot. Additionally, since we wanted to obtain code snippets used in real-world projects, we excluded search results used to solve simple algorithmic problems on platforms, such as LeetCode, which generally involve simple code and may not involve security weaknesses.
@@
We begin by explaining the terminology used in data filtering: the search results under the Repository label are the projects that contain code files, and the search results under the Code label are individual code files. Those code files contain Copilot generated code snippets. In filtering the projects, we followed three rules: (1) for search results under the Repository label, we identified projects that are fully generated by Copilot, as declared in the projects description or the associated README file(s). We retained code files for Python, JavaScript, Java, C++, C#, and Go, which are the main languages supported by Copilot. (2) For search results under the Code label, we retained code files with comments showing the code generated by Copilot. (3) As we mentioned above, we then excluded code files used to solve simple algorithm problems. We provide examples for the three rules in Figs.  3,  4 and  5. As shown in the example in Fig. 3 for the Repository label, we kept all the Python files. In the next example in Fig. 4, we kept the entire file where the Copilot generated code snippet was located. In Fig. 5, the code snippet was removed as it was determined the code just solved a simple algorithmic problem. Meanwhile, for the code files retained under the Repository label, we consider the entire code file as code generated by Copilot. In other words, we assume that all code in the file is generated by Copilot because it was stated in the README file that it was all generated by Copilot. For code files retained under the Code label, we know that the files contain code snippets, perhaps even just a few lines of code, generated by Copilot. Instead of identifying the specific Copilot generated code in this step, we combine the warning messages from the security scan and the code comments in the file to determine whether Copilot generates the code snippet with the security problem (this process is explained further in Section 3.3.2).
After completing the pilot data labeling, the first author checked the rest of the search results, and obtained a total of 465 code snippets. After removing duplicate results, we finally obtained 435 different code snippets. Among them, 249 are from the Repository label, and 186 are from the Code label. Table 3 shows the types and numbers of code snippets obtained.
@@
Figure 3. Example of rule 1: project is fully written by Copilot
@@
Figure 4. Example of rule 2: files with comments showing the code generated by Copilot
@@
Figure 5. Example of rule 3: files used to solve simple algorithm problems
@@
L1 Python 132 119 251
L2 JavaScript 51 28 79
L3 Java 25 18 43
L4 C++ 14 12 26
L6 Go 19 1 20
L5 C# 8 8 16
Total 249 186 435
3.3. Data Pre-processing and Analysis
3.3.1. Data Pre-process
Step 3. CodeQL is a scalable static security analysis tool that is widely used in practice, and enables users to analyze code and detect relevant weaknesses using predefined queries and test suites and supports for multiple languages (including Java, JavaScript, C++, C#, Go, and Python (GitHub, 2021a)). Before using CodeQL to scan the identified code snippets for security weaknesses, we needed to create a CodeQL database for the source code. For interpreted languages like Python and JavaScript, the source code can be directly analyzed, while for compiled languages such as Java, the source code will need to be compiled first and then imported into the CodeQL database. Therefore, we first compiled the code snippets of all compiled languages (i.e., C#, Java, C++, and Go). We removed any code snippets that could not be compiled. For successfully compiled files, we generated the CodeQL database required for queries. At the same time, for interpreted languages Python and JavaScript files, we stored 20 files in each database to improve efficiency, because if we generate a database for an exceptionally large number of files, this would increase the database compilation and scanning time, which is much longer than partitioning them into small databases. In total, we obtained 80 code databases available for CodeQL scanning. Table 4 shows the types and numbers of files in each database.
Table 4. Databases for CodeQL scanning
# Language # Databases: Repository # Databases: Code
L1 Python 7 6
L2 JavaScript 3 2
L3 Java 8 18
L4 C++ 5 12
L5 C# 3 7
L6 Go 13 1
Total 39 46
@@ 3.3.2. Data Analysis
Step 4. We used well-known automated static analysis tools listed by OWASP (OWASP, [n. d.]) to scan the collected code snippets. Since different static analysis tools may use different algorithms and rules to detect security weaknesses, using multiple tools can increase our chances of discovering security issues in the code. To improve the coverage and accuracy of the results, we used two static analysis tools for security checks on each code snippet (i.e., CodeQL plus a dedicated tool for the specific language).
@@
We first used CodeQL to analyze the code in our dataset. The default query suite for the standard CodeQL query package is codeql-suites/<lang>-code-scanning.qls. There are several useful query suites in the codeql-suite directory of each package. For example, the codeql/cpp-queries package contains the following query suites (GitHub, 2021b):
@@
  • Cpp-code-scanning.qls, which is the standard code scanning query for C++. It covers various features and syntax of C++ and aims to discover some common weaknesses in the code.
@@
  • Cpp-security-extended.qls, which includes some more advanced queries than cpp-code-scanning.qls and can detect more security weaknesses.
@@
  • Cpp-security-and-quality.qls, which combines queries related to security and quality, covering various aspects of C++ development from basic code structure and naming conventions to advanced security and performance weaknesses. It aims to help developers improve the security and quality of their code.
In this study, we scanned code snippets using the <language>-security-and-quality.qls test suite related to security weaknesses. These test suites check for multiple security properties and cover many CWEs. For example, the python-security-and-quality.qls test suite for Python provides 168 security checks, the JavaScript test suite provides 203 security checks, and the C++ test suite provides 163 security checks. As the query reports only provide the name and description of the security issues, we manually matched the results in the query reports with the corresponding CWE IDs.
We then selected other popular static security analysis tools for files in each program languages we analyzed. We used the following popular security analysis tools: Bandit for Python, ESLint for JavaScript, Cppcheck for C ++, Findbugs for Java, Roslyn for C# and Gosec for Go. In cases where we could not directly obtain the CWE ID related to the security issue from the scan results, we manually mapped the security attributes to the corresponding CWE for later analysis. We explain the specific correspondences in detail in Section 4.2.
Step 5. We scanned code snippets from the Repository and Code labels, and we filtered the scan results before analyzing them. We first removed the scan results that were repeatedly prompted by two of the tools, then removed the results that were unrelated to the security issue, and finally confirmed that the Copilot generated code indeed caused the results related to the security issue. As we explained in Section 3.2.2, we considered the code snippet from the Repository label to be the entire code file. Therefore, we kept the entire scan results from the Repository label and counted all the security issues they suggested. For the code snippet obtained from the Code label, we started by scanning the code file. If the static analysis tool found a security issue in the code, we located the code snippet in the file according to the line number of the security issue indicated by the scanning result. We determined whether it was generated by Copilot based on the comments before and after the code snippet. If Copilot indeed generated the code snippet with a security issue, we kept the scan result for our subsequent statistics. We further analyzed the filtered scan results in Step 6, detailed in Section 4 according to the specific RQs. We provide our full dataset (including code snippets, full scan results, and filtered results) in our replication package (Fu et al., 2023).

### v2

@@ 3.1. Research Goal and Questions
This study aims to understand the potential security weaknesses in Copilot-generated code produced in real-world GitHub projects. We first collect code snippets generated by Copilot from GitHub projects as our data source, which is the largest source of open-source projects. It should be noted that it is not possible to access all the code generated by Copilot in GitHub projects, as there is no direct way to identify if part of a file was generated by Copilot (i.e., source files do not contain any signatures to indicate if Copilot generates the code). However, we can identify many code snippets by searching the repository description and the comments provided in the code (see the details in Section 3.2.2).
We chose to focus on Python and JavaScript code snippets as they are currently the two most popular programming languages used by developers according to a recent survey in 2022 by GitHub (oct, [n. d.]), and also the ones most frequently used with GitHub Copilot (Zhang et al., 2023). We first analyzed the functionality and application domains of the collected code snippets to get the demographic information of the Copilot-generated code. Next, we performed a security analysis on the identified code. We selected static analysis because of its coverage and the ability to analyze programs without executing them. Dynamic analysis is more sound and precise (as it can reason about the program behaviour) but also lacks coverage (Sui et al., 2020). Using static analysis will allow us to run the analysis on a segment of the program (smaller snippet) without needing the full program to be executed. Static security analysis tools have been widely used by developers and companies (Majdinasab et al., 2024; Ernst, 2003; Pearce et al., 2022). After obtaining the scan results, we manually checked the results to remove false positives reported by the static analysis tools. We finally used CWEs to classify the filtered results for further analysis to answer the RQs.
We conducted this empirical study by following the guidelines of Easterbrook et al. (Easterbrook et al., 2008). The RQs, their rationale, and the research process of this study (see Figure 1) are detailed in the subsections below.
@@
Rationale: Copilot may produce code suggestions that developers accept, but these suggestions may include security weaknesses that could make the program vulnerable. The answer to RQ1 helps understand the frequency of security weaknesses developers encounter when using Copilot in production.
@@
Rationale: Copilot-generated code may contain a variety of security weaknesses (Pearce et al., 2022), and developers should conduct a rigorous security review before accepting the generated code suggestions. Copilot’s documentation notes the following: “users of Copilot are responsible for ensuring the security and quality of their code (GitHub, 2023b)”. By identifying common or recurring weaknesses, developers can be more prepared to prevent, mitigate, or fix these security weaknesses. The answer to RQ2 can help developers better understand possible security weaknesses in the code generated by Copilot, allowing them to be vigilant about the generated code before it is integrated into their programs.
RQ3. How many security weaknesses belong to the MITRE CWE Top-25 security weaknesses?
Rationale: The MITRE list contains the Top 25 most dangerous security weaknesses (The MITRE Corporation, 2023a), offering a benchmark for gauging the severity of weaknesses in the Copilot-generated code (Pearce et al., 2022). The answer to RQ3 can help to understand whether the code generated by Copilot contains widely recognized types of security weakness and Copilot’s ability to handle these recent and common weaknesses.
@@ 3.2. Data Collection and Filtering
We chose GitHub as the primary data source for answering our RQs. GitHub is widely used by developers. As the world’s largest code hosting platform, GitHub contains millions of public code repositories and offers access to a large number of code resources, allowing us to cover multiple programming languages and project types in our study (Cosentino et al., 2017). We identify code snippets generated by Copilot from GitHub projects.
@@ 3.2.1. Code Snippets Collection
Step 1. To identify the code snippets generated by Copilot, we first conducted a pilot search to formulate our search keywords. First, we used “GitHub Copilot” and “Copilot” as our search keywords. As expected, we found that the term “Copilot” is a vague term, referring not only to the code generation tool launched by GitHub, but also to tools in the aviation or telemetry fields. Therefore, using the keyword solely may return irrelevant content that may not be related to the use of the tool. On the other hand, using “GitHub Copilot” can exclude content unrelated to the code generation tool Copilot and narrow the search scope, which is what we have used to locate the code snippets.
However, even with this basic search keyword, we still need to carefully filter the search results to ensure that they are truly related to GitHub Copilot. Although using “GitHub Copilot” increases the relevance of the results to Copilot, these results are not necessarily the code snippets generated by Copilot. It should be noted that many code snippets containing the “GitHub Copilot” keyword in the search results display GitHub Copilot as text. Developers may use them to describe their experience using Copilot to generate code or showcase information related to Copilot. These code snippets are not what we seek as they do not directly relate to the code generated by Copilot. Our target is code generated by Copilot, not code snippets containing the keyword “Copilot”.
Our observations from the pilot search showed that using keywords such as “by GitHub Copilot”, “use GitHub Copilot”, and “with GitHub Copilot” can improve the accuracy of search results. These keywords enable us to focus more on the code generated using Copilot rather than code snippets that contain other content related to Copilot. In addition, since Python and JavaScript are the two most popular programming languages used by developers, and also the most frequently used with GitHub Copilot (see Section 3.1), we further limited the types of code snippets during the search to Python and JavaScript. We collected the search results under the Code label. Considering that some projects declare using GitHub Copilot generated code in their README files or project description provided in GitHub, we decided to retain the results from the Repository label in the search results. Figure 2 shows an example of our search process.
Table 1 reports the search terms we used and the number of search results obtained from GitHub. In this step, we collected a total of 8,157 results, of which 7,749 were from the Code label, and 408 were from the Repository label. The same search result may contain multiple keywords, which means that there are duplicate projects in the collected data. After removing duplicate projects, we obtained a total of 2023 search results, of which 1847 were from the Code label, and 176 were from the Repository label.
@@
1 “By GitHub Copilot” 2549 94
2 “Use GitHub Copilot” 1822 123
3 “With GitHub Copilot” 3378 191
Total 7749 408
@@ 3.2.2. Filtering Code Snippets
Step 2. After obtaining the data from the keyword searches, we further filtered them by not only considering the accuracy of the keywords but also by investigating the project’s documentation, code comments, and other metadata in the search results to determine whether they were generated by GitHub Copilot. Additionally, since we aimed to obtain code snippets used in real-world projects, we excluded search results used to solve simple programming practice problems on platforms, such as bisection lookup and fast sorting tasks from LeetCode. We consider these programming exercises to be more closely related to correctness than security.
@@
We first conducted a pilot data filtration to better filter the search results. We begin by explaining the terminology used in data filtering: the search results under the Repository label are the projects that contain code files, and the search results under the Code label are individual code files. Those code files contain code snippets generated by Copilot. In filtering the projects, we followed three criteria including two inclusion criteria and one exclusion criterion. Inclusion Criterion 1: for search results under the Repository label, we identified projects that are fully generated by Copilot, as declared in the project description or the associated README file(s). We retained code files for Python and JavaScript. Inclusion Criterion 2: For search results under the Code label, we retained code files with comments showing the code generated by Copilot. Exclusion Criterion 1: We excluded code files used to solve simple programming practice problems. We provide examples for the three criteria in Figure 3, Figure 4, and Figure 5. As shown in the example in Figure 3 for the Repository label, we kept all the Python files. In the next example in Figure 4, we kept the entire file where the Copilot-generated code snippet was located. In Figure 5, the code snippet was removed as it was determined that the code just solved a simple algorithmic problem.
Specifically, the pilot data filtering process consisted of the following steps: (1) The first author randomly selected 40 projects and 400 code files from the search results from the Repository label and the Code label, respectively. (2) Two authors independently labelled whether Copilot generated these projects and code files. (3) The first author compared the labelling results by the two coders and calculated the level of agreement between them using the Cohen’s Kappa coefficient (Cohen, 1960). (4) If there were any results they were unsure of or disagreed with, the two authors discussed until they reached an agreement. The Cohen’s Kappa coefficients were 0.79 for the projects (from the Repository label) and 0.85 for the code files (from the Code label), which were both higher than 0.7, indicating a high level of agreement between the two coders and ensuring a good accuracy of the labelling results. After completing the pilot data labelling, the first author checked the rest of the search results (1,447 code files and 136 projects). After removing the duplicate results, we retained 79 projects under the Repository label and 180 code files under the Code label. Step 3. To further get the application domains represented in the dataset, we categorized the projects based on the project’s description documents and the specific function of the code. Projects were classified into one of the following categories: Games, Web Applications, Utility Tools, AI Applications, Network Communication Applications, and Others. Figure 6 shows the classification and distribution of the application domains of the Copilot-generated code in the dataset from the Repository and Code labels respectively.
For the code files retained under the Repository label, we consider the entire code file as code generated by Copilot. In other words, we assume that Copilot generates all the code in the file because it was stated in the README file that it was all generated by Copilot. For code files retained under the Code label, we know that the files contain code snippets, perhaps even just a few lines of code, generated by Copilot. Instead of identifying the specific Copilot-generated code in this step, we combine the warning messages from the security scan and the code comments in the file to determine whether Copilot generates the code snippet with the security problem (this process is explained further in Section 3.3.2). As a result, we obtained 452 code files with different contents, 272 from the Repository label and 180 from the Code label. Table 2 gives the type and number of code files. A curated dataset, comprising all the data collected during the research process, has been made available (Fu et al., 2024).
@@
Figure 3. Example of Inclusion Criterion 1: projects fully written by Copilot
@@
Figure 4. Example of Inclusion Criterion 2: files with comments showing the code generated by Copilot
@@
Figure 5. Example of Exclusion Criterion 1: files used to solve simple algorithm problems
(a) Code from the Repository label
(b) Code from the Code label
Figure 6. Application domains of the Copilot-generated code in our dataset
@@
L1 Python 143 134 277
L2 JavaScript 129 46 175
Total 272 180 452
3.3. Data Processing and Analysis
3.3.1. Data Processing
Step 4. CodeQL is a scalable static security analysis tool that is widely used in practice and allows users to analyze code and detect relevant weaknesses using predefined queries and test suites and supports for multiple languages (including Java, JavaScript, C++, C#, Go, and Python (GitHub, 2023a)). Before using CodeQL to scan the identified code snippets for security weaknesses, we needed to create a CodeQL database for the source code. Source code can be directly analyzed for interpreted languages like Python and JavaScript without being compiled into intermediate code. To speed up the scan, we stored 20 files in each database using the command line to generate the database needed for CodeqQL queries. It should be noted that generating a database for an exceptionally large number of files would increase the database compilation and scanning time, which is much longer than partitioning them into small databases. In total, we obtained 25 code databases available for CodeQL scanning.
@@ 3.3.2. Data Analysis
Step 5. We used well-known automated static analysis tools listed by OWASP (OWASP, [n. d.]) to scan the collected code snippets. Static analysis has been widely used to detect security issues in code, for its ability to analyze programs without execution (Dunlap et al., 2023). While dynamic analysis offers more precise insights by reasoning about program behavior, it suffers from limited coverage (Sui et al., 2020). Employing static analysis enables us to analyze program segments (smaller snippets) without requiring the entire program to be executed. Since different static analysis tools may use different algorithms and rules to detect security weaknesses, using multiple tools can increase our chances of discovering security issues in the code. To improve the coverage and accuracy of the results, we used two static analysis tools for security checks on each code snippet (i.e., CodeQL plus one dedicated tool for the specific language, Bandit for Python and ESLint for JavaScript).
@@
We first used CodeQL to analyze the code in our dataset. The default query suite for the standard CodeQL query package is codeql-suites/<lang>-code-scanning.qls. Each package has several query suites in the codeql-suite directory. For example, the codeql/python-queries package contains the following query suites (GitHub, 2023):
@@
  • Python-code-scanning.qls, the standard scanning query for Python. It covers various features and syntax of Python and aims to discover some common weaknesses in the code.
@@
  • Python-security-extended.qls, which includes some more advanced queries than Python-code-scanning.qls and can detect more security weaknesses.
@@
  • Python-security-and-quality.qls, which combines queries related to security and quality, covering various aspects of Python development, from basic code structure and naming conventions to advanced security and performance weaknesses. It aims to help developers improve the security and quality of their code.
In this study, we scanned code snippets using the <language>-security-and-quality.qls test suite. These test suites check for multiple security properties and cover many CWEs. For example, the python-security-and-quality.qls test suite for Python provides 168 security checks, and the JavaScript test suite provides 203 security checks. As the query reports only provide the name and description of the security issues, we manually matched the results in the query reports with the corresponding CWE IDs. We then selected one additional static security analysis tool for files in each programming language we analyzed: Bandit for Python and ESLint for JavaScript. As explained in Section 3.2, we considered the code snippet from the Repository label to be the entire code file, while the code snippet from the Code label exists in the code file, with the exact number of lines unspecified at this stage. Step 6. We scanned code snippets from the Repository and Code labels, and we filtered the scan results before analyzing them. In this step, we adopted part of the strategies used by developers when they perform different tasks with static analysis tools: warning prioritization and determining whether a warning is a false positive or a true report (Do et al., 2020). We first performed an initial filtering of the results based on the priority of the warnings. Specifically we removed the repeated scan results that were reported by the two tools, then removed the scan results that were not security issues, such as recommendations, to get the initial scan results. Subsequently, to confirm that the security issues were actually caused by the Copilot-generated code, the first author manually checked the scan results. For the initial results obtained under the Repository label, we checked them one by one. Specifically, we determined whether the corresponding location of the code snippet indeed had a security issue based on the line number information provided by the scan results. For the initial results obtained under the Code label, we confirmed whether the security issues truly existed and verified if the security issues were caused by Copilot-generated code. Specifically, after scanning the code file, we pinpointed the code snippet within the file based on the line number of the security issue indicated in the scan results. We assessed whether it was generated by Copilot by checking the surrounding comments and determined whether a security issue existed in that particular context. We further analyzed the filtered scan results in Step 7, detailed in Section 4, according to the specific RQs.
We provide the complete dataset (including code snippets, full scan results, and filtered results) in our replication package (Fu et al., 2024).